OpenStudy (anonymous):

Which expression is equivalent to the expression 4 1/8 divided by 5 2/5 33/8 x 27/5 33/8 x 5/27 8/41 x 52/5 41/8 x 5/52

OpenStudy (anonymous):

help meh out plz

OpenStudy (skullpatrol):

What grade?

OpenStudy (skullpatrol):

@pandaluvs are you still there?

OpenStudy (anonymous):

oh i am sorry please, i am here,,,, i am in 6th grade i am so sowwy,

OpenStudy (anonymous):

middle school

OpenStudy (anonymous):

do u look like your profile picture in real life???? (just asking)

OpenStudy (anonymous):

umm if you dont mind, would u type faster???

OpenStudy (skullpatrol):

Ok, the first thing you need to learn is how to convert a number in the "mixed" form like $$\Huge 4\frac{1}{8}$$ into a proper fraction form like $$\Huge \frac{33}{8}$$ can you do that?

OpenStudy (anonymous):

yesh i can

OpenStudy (anonymous):

do u know that i have a panda?

OpenStudy (anonymous):

4 real, i adopted him when he was a newborn

OpenStudy (anonymous):

we live in south carolina in a ranch

OpenStudy (skullpatrol):

Convert $$\Huge 5\frac{2}{5}$$

OpenStudy (anonymous):

dis is him last week,

OpenStudy (anonymous):

ok i will convert

OpenStudy (anonymous):

hold on just a sec

OpenStudy (anonymous):

27/5

OpenStudy (anonymous):

\[27/5\]

OpenStudy (anonymous):

dude!!!!!! r u there

OpenStudy (skullpatrol):

Now we need to use the rule of division for fractions.

OpenStudy (anonymous):

did u read my panda thing????????

OpenStudy (anonymous):

do u know that i have a panda? 4 real, i adopted him when he was a newborn we live in south carolina in a ranch

OpenStudy (skullpatrol):

$$\Huge \frac{a}{b} \div \frac{c}{d}= \frac{a}{b}*\frac{d}{c}$$

OpenStudy (anonymous):

and he lives with us

OpenStudy (anonymous):

what the HECK does tht mean????

OpenStudy (anonymous):

ooh yeah the rule of fractions... =p

OpenStudy (skullpatrol):

To divide by a fraction, you multiply by its reciprocal.

OpenStudy (anonymous):

dude u take sooo long to type, like for example, what u just wrote i could have written it under 05 seconds!!!

OpenStudy (skullpatrol):

Sorry, I am old :(

OpenStudy (anonymous):

puh leez its like u never end!!!!

OpenStudy (anonymous):

how old r u?

OpenStudy (anonymous):

cause if ur old then ur profile pic lies

OpenStudy (anonymous):

dude seriously, how old r u???

OpenStudy (anonymous):

just type a NUMBER!!!!!

OpenStudy (skullpatrol):

Probably older than your teacher :-) Have you learned the "Division Rule for Fractions" in school?

OpenStudy (anonymous):

yes i have, but my question is not being answered, HOW OLD ARE YOU?????? i bet you 50.000.000 bucks ur not older than my dad!!!

OpenStudy (skullpatrol):

Why do you use decimal points "." instead of commas"," ???

OpenStudy (anonymous):

cause im lazy now will u pleez answer my friking question????

OpenStudy (anonymous):

x_x ANSWER DA' QUESTION x_x

OpenStudy (skullpatrol):

Which expression is equivalent to the expression 4 1/8 divided by 5 2/5 33/8 x 27/5 33/8 x 5/27 8/41 x 52/5 41/8 x 5/5

OpenStudy (skullpatrol):

can you answer your math question now?

OpenStudy (anonymous):

yes, now i can

OpenStudy (skullpatrol):

That would not be appropriate.

OpenStudy (anonymous):

@hero @hero @hero @hero @ hero

OpenStudy (skullpatrol):

Do you want to talk about math?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

please make this faster, it has been 40 minutes!

OpenStudy (skullpatrol):

Why can you never divide by zero?

OpenStudy (anonymous):

because u will get 0.0000000000000007 for example

OpenStudy (anonymous):

and u do not divide it is like doing nothing!!!

OpenStudy (skullpatrol):

0.0000000000000007 is a number just like any other so that reason is not right

OpenStudy (anonymous):

im not that good at giving examples ok???

OpenStudy (anonymous):

STOP QUESTIONING AND GET TO THE POINT PLEASE

OpenStudy (anonymous):

i REFUSE TO BE HERE ALL DAY!!!

OpenStudy (anonymous):

YOU KNOW WHAT? IF UNDER 50 SECONDS WE DONT GET TO A FINAL ANSWER I WILL PUT IN A RANDOM ONE, CAUSE THIS IS TIMED!!!

OpenStudy (skullpatrol):

Dividing by zero would mean multiplying by the reciprocal of zero.

OpenStudy (anonymous):

50

OpenStudy (anonymous):

49

OpenStudy (anonymous):

48

OpenStudy (anonymous):

47

OpenStudy (anonymous):

46

OpenStudy (anonymous):

HURRY! 45

OpenStudy (anonymous):

44

OpenStudy (anonymous):

43

OpenStudy (anonymous):

42

OpenStudy (anonymous):

41

OpenStudy (anonymous):

40

OpenStudy (skullpatrol):

Hello??? you are not paying attention, look above: Dividing by zero would mean multiplying by the reciprocal of zero.

OpenStudy (anonymous):

IF U DONT HURRY 39 WILL GO BY!!!! 38

OpenStudy (anonymous):

37

OpenStudy (anonymous):

THIS IS A TIMED TEST!!! 36 35

OpenStudy (skullpatrol):

That is what the Division Rule for Fractions says, right?

OpenStudy (anonymous):

YES 34 33 32

OpenStudy (anonymous):

31 30

OpenStudy (anonymous):

29 !!! 28 !!!

OpenStudy (anonymous):

27

OpenStudy (anonymous):

O MI GOOODDD!!!!!!!!!!

OpenStudy (anonymous):

26 25

OpenStudy (anonymous):

24 23

OpenStudy (anonymous):

22 21

OpenStudy (skullpatrol):

troll

OpenStudy (anonymous):

IS IT A,B,C OR D!!!!!!!1

OpenStudy (skullpatrol):

troll

OpenStudy (anonymous):

17 16 15

OpenStudy (skullpatrol):

troll

OpenStudy (anonymous):

WHY R U DOING THAT I REPORTED U TO THE MODERATORS ALREADY !!!

OpenStudy (anonymous):

10 9

OpenStudy (skullpatrol):

i pwn u troll

OpenStudy (anonymous):

WHAT !?!?!

OpenStudy (anonymous):

OK IT IS OVER, you are the worst! i will NEVER!! recomend u to nobody!!!!!!!!!!

OpenStudy (anonymous):

i am not giving u a medal!

OpenStudy (skullpatrol):

Dividing by zero would mean multiplying by the reciprocal of zero. But zero has no reciprocal because zero times any number is zero, not 1. Therefore division by zero has no meaning in the set of real numbers.

OpenStudy (anonymous):

does that help??? NOOO!!!!!!!

OpenStudy (anonymous):

what ever, u wasted my time and thanks to you i have a new f in math!!!!!

OpenStudy (skullpatrol):

Try to think about it. It is the hardest thing to learn and understand in math.

OpenStudy (anonymous):

an F!!!!